Use the following information to answer the next three questions.DO NOT USE EXCEL PLEASE. Formula to be used :

p= forward premium: (F-S)/S. (direct quotes)

•S=current spot rate

0 0
Answer #1

i) Spot rate (¥ per C$) = 1 / (C$ per $ * $ per ¥)

Spot rate (¥ per C$) = 1 / (C$1.25 * $0.008)

Spot rate (¥ per C$) = 1 / C$0.01

Spot rate (¥ per C$) = ¥100/C$

Now, we will calculate forward rate as per IRP (Interest rate parity) theory,

F / S = (1 + i d) / (1 + i f)

Here,

F = Forward rate, S = Spot rate @ ¥100

i d = Interest rate in domestic country ie Japan @ 7% or 0.07

i f = Interest rate in foreign country ie. Canada @ 2% or 0.02

We will put the values into the formula,

F / ¥100 = (1 + 0.07) / (1 + 0.02)

F / ¥100 = 1.07 / 1.02

F / ¥100 = 1.0490

F = 1.0490 * ¥100

F = ¥104.90

ii) Spot rate (C$ per ¥) = 0.01

Forward rate (C$ per ¥) given = 1 / ¥ per C$

Forward rate (C$ per ¥) = 1 / ¥106.50 = C$0.0094

Now,

Discount/premium formula = (F - S) / S

Here,  

F = Forward rate @ C$0.0094

S = Spot rate @ C$0.01

Put the values into the formula,

Discount or premium on ¥ = (0.0094 - 0.01) / 0.01

Discount or premium on ¥ = - 0.06

¥ currency is at discount against C$ as it is having negative value.

iii) Covered interest arbitrage:

Step 1: Quoted forward rate @ ¥106.50 > calculated forward rate @ ¥104.90. Hence borrow in Japan @7% and invest in Canada @ 2%

Borrowed ¥100 @ 7% and total payable = ¥100 + 7% = ¥107

Step 2: Convert ¥100 into C$ @ spot rate = ¥100 ie C$1

Step 3: Invest C$ 1 @ 2% for one year and recievable after a year = 1 + 2% = C$1.02

Step 4: Convert C$1.02 into ¥ @ forward rate = ¥106.50 ie. C$1.02 * ¥106.50 = ¥108.63

Step 5: Net benefit = Receivable (step 4) - payable (step 1) = ¥108.63 - ¥107 = ¥1.63

Now,

Return % = Gain / spot rate * 100

Return % = 1.63 / 100 * 100 = 1.63%